OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = 1/2x - 8. A) y = 1/2x + 7 B) y = 1/2x + 5 C) y = 2x + 7 D) y = 2x + 5 17.

OpenStudy (anonymous):

the 17 at the bottom shouldent be there oops

OpenStudy (nali):

for this quesiton just use the following equation Y-y = m (X-x) for the small letters x and y just plug in the point that u have and for the slope use the same slope, which is 1/2 since the lines are parallel

OpenStudy (anonymous):

im still not sure i understand /:

OpenStudy (anonymous):

y-6=1/2(x+2) 2y-12=x+2 2y=x+14

OpenStudy (nali):

Y-y= m (X-x) look it is all just depends on the plugging in of numbers what seems confusing to u ?

OpenStudy (anonymous):

What @Nali is trying to say. \[y - y_1 = m(x - x_1)\]where m = 1/2 \(x_1 = -2\) \(y_1 = 6\)